OpenStudy (anonymous):

Find the quotient and remainder of the division problem. x^3+1 ------ x+3

OpenStudy (anonymous):

(x^3) + 1 or x^(3+1) ?

OpenStudy (anonymous):

\[\frac{ x ^{3}+1 }{ x+3 }\] @ryclark

OpenStudy (anonymous):

x^3 + 1 = (9 - 3x + x^2) * (3 + x) - 26 ; quotient: (9-3x+x^2), remainder: -26